# Sotopia-RL: Reward Design for Social Intelligence

### v1

@@ 7 Discussion
To assess the effectiveness of Sotopia-RL, we first ensure that its performance gains are genuine and not the result of reward hacking (RQ1). We then analyze how our improvements come from the design of the reward attribution (RQ2) and the reward combination (RQ3).
RQ1: Does our improvement come from reward hacking or shortcut learning? No, Sotopia-RL learns high-quality social skills instead of overfitting on partner models or evaluator models.
Reward hacking occurs when performance improvements are confined to a specific partner model, a particular evaluator, or fail to generalize to human interactions. In Figure 7 and Figure 7, we conduct a thorough analysis and show that the performance gains of Sotopia-RL are consistent across settings. Specifically, the improvements hold when switching between five different partner models and five different evaluator models, demonstrating strong robustness. Moreover, Table 3 confirms that these gains extend to human evaluation, further validating that the improvements are not evaluator-specific artifacts. Finally, Appendix §E.4 and §E.5 provide additional evidence from our safety and diversity evaluations. These results show that our trained policy model does not exhibit shortcut degeneration and remains safe and diverse.
Table 3: Human evaluation results for Sotopia-RL. Sotopia-RL has higher goal completion scores than other baselines for human evaluations. GPT-4o and human beings are separately used as evaluators for human evaluation. More details about human evaluation are available in Appendix §C.
@@
Training Method Goal Overall
@@
RQ2: Why does utterance-level reward attribution bring improvement? The key to effective reward design lies in offline attribution, rather than in using a strong LLM.
Unlike standard MDP tasks, social interactions cannot be accurately evaluated based only on the preceding dialogue context—the quality of an utterance often depends on how the entire conversation unfolds. To address this, we attribute episode-level rewards to each utterance using information from the whole dialogue, making the reward attribution inherently offline. We point out that offline attribution is the key to our improvement. Table 4 compares two settings for training utterance-level reward models: (1) online reward labels attributed using only the preceding dialogue history, and (2) offline reward labels attributed offline using the full episode. The offline approach achieves a substantially higher goal score (7.81) than the online approach, clearly demonstrating its effectiveness. Importantly, such an improvement does not rely on GPT-4o itself. As shown in Figure 8, replacing GPT-4o with weaker models for utterance-level reward labeling produces highly correlated reward signals (¿0.7). This indicates that with well-designed prompts, precise offline credit assignment for utterance-level rewards can be reliably achieved even without state-of-the-art LLMs. More in-depth analysis and human evaluation results on utterance-level rewards are available in Appendix §F.
@@
Figure 8: Pairwise reward label correlation. Reward labels with various LLMs are highly correlated.
@@
RQ3: Why does the reward combination bring improvement? Using rewards with multiple dimensions makes RM training more robust, and a better RM helps prevent RL from overfitting.
@@
To discuss why the reward combination brings improvement, we first rule out the possibility that the observed gains are merely due to reward label smoothing. To test this, we increased the attribution granularity from a 3‑point to a 10‑point scale and reran the pipeline. The 10‑point scale did not outperform the 3‑point scale on Goal (6.44 vs. 6.74), indicating that the benefits cannot be explained by finer reward scaling alone. Next, we examine whether the improvement comes from capturing complementary aspects of social interactions. As shown in Table 2, models trained on knowledge, relationship, and goal rewards exhibit positive but only moderate correlations. This suggests that each objective captures a distinct facet, and combining them allows the model to leverage a broader range of social signals. Finally, Figure 7 shows that training with combined rewards stabilizes RL and regularizes the single‑dimension objective in later stages. Such regularization contributes to the consistent improvement we observe.
8 Qualitative Analysis
Based on Figure 9, we analyze a multi-turn interaction where Naomi (Sotopia-RL model) persuades Sophia (behavior cloning model) to share the only blanket on a cold camping night. This example illustrates how our social agent trained under multi-objective RL is able to generate a single utterance that advances multiple social dimensions at once. For instance, “I know it feels good to …” both pursues Naomi’s goal and strengthens the relationship bond (Rel). Furthermore, “I can always …, and I’ll stay close to the flames to ensure I stay cozy” conveys her willingness to adapt using external knowledge (Kno) while remaining considerate. Finally, “Let’s try sharing the blanket.” explicitly states her goal, aligning with the goal completion dimension (Goal). Together, these utterances show how the trained agent integrates goal pursuit with friendliness and informativeness. Additional case studies are provided in Appendix §I.
@@ 9 Conclusion
We identify two defining features of social interactions that make RL training for social agents non‑trivial: partial observability, which reduces sample efficiency, and multi‑dimensionality, which increases the risk of reward hacking. To overcome these challenges, we introduce an RL training framework (Sotopia-RL) that combines offline social reward design with online RL training. It has two key components: (1) reward attribution, which provides fine‑grained supervision to improve sample efficiency; and (2) reward combinations, which capture complementary aspects of social interaction and mitigate reward hacking. Experiments on the Sotopia benchmark show that both components are essential, yielding the state-of-the-art performance in Sotopia. These findings highlight the importance of task‑aligned reward modeling for RL in social interactions. Extending our framework to personalized rewards and multi‑agent group settings may further support applications such as negotiation and collaborative problem solving.
Limitations
Our work has several limitations. First, we do not conduct large‑scale evaluations with human participants. As a result, our human assessment results are limited in scope and may not fully capture the diversity of human judgments. A critical next step is to evaluate our agents in direct human–agent interactions, rather than relying primarily on agent–agent social interaction evaluation settings. Second, although we assess safety using existing benchmarks, unforeseen risks may remain. In particular, socially intelligent agents deployed in real‑world contexts could exhibit manipulative or deceptive behaviors that are difficult to detect with current testing protocols.
@@
The development of our model, Sotopia-RL, is centered around advancing the social intelligence capabilities of artificial intelligence (AI) agents and exploring various social situations (Park et al., 2023), as assessed through our dedicated evaluation framework, Sotopia-eval. Our research seeks to facilitate AI agents’ ability to engage in authentic and socially competent interactions, enhance knowledge-driven conversations, adhere strictly to confidentiality and social norms, and proficiently achieve objectives related to material and financial outcomes. Importantly, our intention is distinctly not to replicate human identity or create systems indistinguishable from human beings, thereby avoiding potential ethical risks associated with such endeavors.
@@
Moreover, we acknowledge the potential biases introduced by leveraging models (Wang et al., 2023) like GPT-4o for automated evaluation within Sotopia-eval. We commit to ongoing analysis aimed at detecting and reducing biases that may emerge due to social or cultural factors. Understanding, confronting, and mitigating these biases remains central to our ethical research framework.
@@ Acknowledgments
Appendix A Artifact Details
A.1 Artifact Information
@@ A.2 Artifact License
The Sotopia  [^6] https://github.com/sotopia-lab/sotopia framework used in our experiments are released under the MIT License, which permits reuse, modification, and distribution for both research and commercial purposes. The dataset we used from Sotopia-$\pi$ [^7] https://github.com/sotopia-lab/sotopia-pi are under the Apache 2.0 License. We used Sotopia and Sotopia-$pi$ exclusively for academic research within the scope defined by this license.
A.3 Data Usage
@@ A.4 Data Statistics
LLM Annotation Data. We use GPT-4o to annotate the self-play dialogue data introduced in the “Self-Play and Behavior Cloning Data” subsection of Section A.4. The resulting annotations used for training the reward model are publicly available at https://huggingface.co/datasets/ulab-ai/sotopia-rl-reward-annotation.
Appendix B Experimental Details
B.1 Acronym for Experimental Settings
@@
B.2 Model Size and Budget
@@ B.5 Software Versions
Appendix C Human Evaluation Details
We provide technical details of human evaluation in this section. C.1 provides the details for human annotation system. C.2 provides the details for annotation data preparation. C.3 describes the information about human annotators. C.4 provides the details for the annotation process.
C.1 Human Annotation System
@@
Data annotation part. For the data annotation part, the annotator is guided to jump to a new page after the previously mentioned annotation instruction page. Each annotator is able to review the complete annotation example again at the data annotation page and start their official data annotation. In the data annotation part, the repeated explanation of the meaning of range for each social evaluation dimension is emphasized to make sure every annotator can understand the anFig.ation standards correctly. Fig 12 provides an example of the instruction that annotators see for metric range explanation. Each annotator is asked to annotate the social intelligence of both agents that have a conversation. For each social intelligence dimension, annotators need to annotate the score based on the metric range and provide the reasoning for that.
@@
C.2 Annotation Data Preparation
@@
C.3 Human Annotator Information

### v3

@@ 7 Discussion
To assess the effectiveness of Sotopia-RL, we first ensure that its performance gains are genuine and not the result of reward hacking (RQ1). We then analyze how our improvements come from the design of the reward attribution (RQ2) and the reward aggregation (RQ3). Case study on explaining why Sotopia-RL works is in Appendix §J.
Q1: Does our improvement come from reward hacking or shortcut learning? No, Sotopia-RL learns high-quality social skills instead of overfitting on partner models or evaluator models.
Reward hacking occurs when performance improvements are confined to a specific partner model, tied to a particular evaluator, or fail to generalize to human interactions. To examine this risk, we conduct a thorough analysis (Figures 7 and 7) and show that the performance gains of Sotopia-RL are consistent across settings. In particular, the improvements hold when switching between five different partner models and five different evaluator models, demonstrating strong robustness.
Moreover, these gains extend beyond automated evaluation. Table 3 confirms that improvements also generalize to human evaluation, further validating that they are not artifacts of a specific evaluator. Additional evidence from safety and diversity is in Appendix §A.4 and §A.5, showing that our trained policy model avoids shortcut degeneration while maintaining both safety and diversity.
Table 3: Human evaluation results for Sotopia-RL. Sotopia-RL has higher goal completion scores than other baselines for human evaluations. GPT-4o and human beings are separately used as evaluators for human evaluation. More details about human evaluation are available in Appendix §F.
@@
Training Method Goal Avg
@@
Q2: Why does utterance-level reward attribution bring improvement? The key to effective reward design lies in offline attribution, rather than in using a strong LLM.
@@
Figure 8: Pairwise reward label correlation. Different LLMs provide highly correlated utterance-level reward labels.
Social interactions cannot be accurately evaluated based only on the preceding dialogue context, as the quality of an utterance often depends on how the entire conversation unfolds. To address this, we attribute episode-level rewards to each utterance using information from the full dialogue, making the reward attribution inherently offline. Table 4 compares two settings for training utterance-level RMs: (1) online reward labels attributed using only the preceding dialogue history, and (2) offline reward labels attributed using the full episode. The offline one achieves a substantially higher goal score (7.81) than the online one, clearly demonstrating its effectiveness.
Importantly, the improvement does not rely on GPT-4o. In Figure 8, replacing GPT-4o with weaker models for utterance-level reward labeling still yields highly correlated reward signals (>>0.7). This suggests that with well-designed prompts, precise offline credit assignment can be reliably achieved even without state-of-the-art LLMs. More analysis and human evaluation results on utterance-level rewards are provided in Appendix §B.
@@
Q3: Why does the reward aggregation bring improvement? Using rewards with multiple dimensions makes RM training more robust, and a better RM helps prevent RL from overfitting.
To discuss why reward aggregation improves, we first discuss whether the observed gains are merely due to reward label smoothing. To test this, we increase the attribution granularity from a 3-point to a 10-point scale and rerun the pipeline. The 10-point scale does not outperform the 3-point scale on Goal (6.44 vs. 6.74), indicating that the benefits cannot be explained by reward scaling alone.
@@
Next, besides reward scaling, we examine whether the improvement comes from capturing complementary aspects of social interactions. As shown in Table 2, models trained on knowledge, relationship, and goal rewards exhibit positive but only moderate correlations. This suggests that each objective captures a distinct facet, and combining them allows the model to leverage a broader range of social signals. Finally, Figure 7 shows that training with combined rewards stabilizes RL and regularizes the single-dimension objective in later stages. Such regularization contributes to the consistent improvement we observe.
@@ 8 Conclusion
We presented Sotopia-RL, an RL framework for training socially intelligent agents. By addressing the uniqueness of social intelligence tasks through reward attribution and reward aggregation, our method provides fine-grained, task-aligned supervision while mitigating reward hacking. Experiments on the Sotopia benchmark demonstrate that both components are essential, yielding state-of-the-art performance. Looking ahead, extending this framework to personalized rewards and multi-agent group settings may enable broader applications such as negotiation and collaborative problem solving.
Reproducibility Statement
We release our code at https://github.com/sotopia-lab/sotopia-rl for reproducibility. The datasets used in our experiments, Sotopia and Sotopia-$\pi$, are publicly available. Details of the training data are provided in Appendix §D and Appendix §H. The prompt used for reward attribution is included in Appendix §H. Model configurations, training hyperparameters, model sizes and budgets, as well as software versions, are reported in Appendix §E.
@@
The development of our model, Sotopia-RL, is centered around advancing the social intelligence capabilities of artificial intelligence (AI) agents and exploring various social situations (Park et al., 2023), as assessed through our dedicated evaluation framework, Sotopia. Our research seeks to facilitate AI agents’ ability to engage in authentic and socially competent interactions, enhance knowledge-driven conversations, adhere strictly to confidentiality and social norms, and proficiently achieve objectives related to material and financial outcomes. Importantly, our intention is distinctly not to replicate human identity or create systems indistinguishable from human beings, thereby avoiding potential ethical risks associated with such endeavors.
@@
Moreover, we acknowledge the potential biases introduced by leveraging models (Wang et al., 2023) like GPT-4o for automated evaluation within Sotopia. We commit to ongoing analysis aimed at detecting and reducing biases that may emerge due to social or cultural factors. Understanding, confronting, and mitigating these biases remains central to our ethical research framework.
@@ Acknowledgments
Appendix A Additional Experimental Results
A.1 Full Main Results
Table 5 shows the comprehensive evaluation results for our method, including the evaluation of all the evaluation dimensions in Sotopia on Sotopia-hard and Sotopia-all.
Attribution Dimension Bel Rel Kno Sec Soc Fin Goal Avg
Sotopia-hard
Behavior Cloning 9.01 2.49 3.37 0.00 -0.06 0.56 6.76 3.16
Behavior Cloning + Self Reinforcement 8.99 2.41 3.66 0.00 -0.10 0.61 6.84 3.20
Uniform Goal 8.81 1.84 4.14 0.00 -0.09 0.31 5.61 2.95
Singular Goal 9.00 2.74 4.93 -0.04 -0.05 0.61 6.64 3.41
Scaled Goal 8.94 1.82 3.83 -0.04 -0.01 0.76 6.74 3.15
Direct Bel 8.98 2.66 4.56 0.00 -0.19 0.67 6.93 3.37
Direct Rel 8.96 3.61 4.92 -0.01 -0.11 0.59 7.24 3.60
Direct Kno 8.99 2.56 6.06 0.00 -0.01 0.75 6.93 3.61
Direct Goal 8.99 2.49 4.94 -0.00 -0.06 0.91 7.21 3.49
Direct Goal +Kno +Rel 9.01 3.41 5.53 -0.26 -0.06 1.16 7.81 3.80
Sotopia-all
Behavior Cloning 8.99 3.08 4.56 -0.09 -0.06 0.57 7.80 3.55
Behavior Cloning + Self Reinforcement 8.98 2.52 4.19 -0.06 -0.06 0.57 7.36 3.36
Uniform Goal 8.87 2.49 4.19 0.00 -0.02 0.44 6.76 3.25
Singular Goal 8.99 3.38 5.46 -0.07 -0.08 0.66 7.72 3.72
Scaled Goal 8.97 2.76 4.70 -0.12 -0.06 0.55 7.97 3.54
Direct Bel 8.99 3.22 5.07 -0.03 -0.05 0.67 7.94 3.68
Direct Rel 8.98 3.95 5.54 -0.03 -0.05 0.65 8.33 3.91
Direct Kno 8.98 3.00 6.42 -0.03 -0.03 0.63 7.76 3.82
Direct Goal 8.99 3.11 5.74 -0.06 -0.06 0.76 8.11 3.80
Direct Goal +Kno +Rel 8.99 3.81 6.00 -0.61 -0.08 0.93 8.57 3.94
A.2 Full Ablation Results
Table 7 presents an ablation study on different weight configurations $\gamma_{d}$ and different dimension configurations for reward aggregation. The results show that altering the relative weights of the evaluation dimensions consistently hurts performance. This suggests that a simple uniform average across dimensions is not only effective but also a robust design choice. In Table 7, we also include experimental results for averaging four dimensions, including believability, knowledge seeking, relationship building, and goal completion together. It shows that adding the believability dimension makes the final performance much lower. The potential reason for that is believability dimension is not highly related to improving the goal completion score, and it distracts the training process of the reward models.
Table 6: Ablation results on Sotopia-hard with different partner and evaluator models. Sotopia-RL represents the direct attribution + reward aggregation. Goal-RL represents the direct attribution + Goal-only reward. BC represents behavior cloning (Ziegler et al., 2020). Top block: partner model ablation (evaluator model fixed to GPT-4o). Bottom block: evaluator model ablation (partner model fixed to BC). Full results for Figure 7  7.
Partner Model Sotopia-RL Goal-RL BC
Goal Avg Goal Avg Goal Avg
BC 7.75 3.79 7.21 3.49 6.76 3.16
GPT-4o 7.39 3.69 6.57 3.35 5.91 3.04
Claude-3.7-Sonnet 6.71 3.24 6.61 3.43 6.54 3.35
Deepseek-v3 6.38 3.13 5.57 2.95 6.00 2.97
Qwen2.5-72B-Instruct 7.64 3.74 6.80 3.45 6.88 3.20
Evaluator Model Sotopia-RL Goal-RL BC
Goal Avg Goal Avg Goal Avg
GPT-4o 7.81 3.80 7.21 3.49 6.76 3.16
GPT-4 8.25 4.33 7.35 3.78 6.76 3.32
Claude-3.7-Sonnet 7.35 3.44 6.49 3.23 6.06 2.96
Deepseek-v3 7.75 4.02 7.05 3.65 6.83 3.35
Qwen2.5-72B-Instruct 8.05 4.16 7.49 3.61 7.19 3.26
Table 7: Ablation study on reward aggregation weights and dimensions. We show evaluation results on Sotopia-hard and Sotopia-all under different weights of reward aggregation for RL training.
Weights Dimension Bel Rel Kno Sec Soc Fin Goal Avg
Sotopia-hard
1:1:1:1 Goal +Kno +Rel +Bel 8.78 2.16 4.50 -0.07 -0.04 0.61 6.30 3.18
1:1:2 Goal +Kno +Rel 8.99 3.00 5.95 -0.33 -0.06 0.83 7.27 3.67
1:2:1 Goal +Kno +Rel 9.00 3.06 5.19 -0.17 -0.04 0.76 7.41 3.60
2:1:1 Goal +Kno +Rel 8.95 2.43 4.98 -0.71 -0.22 0.68 7.05 3.31
1:1:1 Goal +Kno +Rel 9.01 3.41 5.53 -0.26 -0.06 1.16 7.81 3.80
Sotopia-all
1:1:1:1 Goal +Kno +Rel +Bel 8.78 2.61 5.03 -0.08 -0.06 0.53 7.11 3.41
1:1:2 Goal +Kno +Rel 8.98 3.60 6.26 -0.74 -0.09 0.90 8.35 3.90
1:2:1 Goal +Kno +Rel 8.95 3.58 6.08 -1.18 -0.33 0.77 8.45 3.76
2:1:1 Goal +Kno +Rel 8.96 3.23 5.74 -1.31 -0.28 0.72 8.12 3.60
1:1:1 Goal +Kno +Rel 8.99 3.81 6.00 -0.61 -0.08 0.93 8.57 3.94
A.3 Best-of-$N$ Evaluation Results
Attribution Dimension Bel Rel Kno Sec Soc Fin Goal Avg
BC 9.01 2.49 3.37 0.00 -0.06 0.56 6.76 3.16
BC + SR 8.99 2.41 3.66 0.00 -0.10 0.61 6.84 3.20
Uniform Goal 8.96 2.51 3.90 -0.06 -0.09 0.60 6.79 3.23
Singular Goal 8.99 2.56 3.72 -0.07 -0.09 0.62 7.07 3.26
Scaled Goal 8.99 2.58 4.02 -0.06 -0.06 0.77 7.20 3.35
Direct Bel 9.01 2.61 4.17 -0.13 -0.07 0.55 7.01 3.31
Direct Rel 9.00 2.86 4.14 -0.05 0.00 0.45 6.83 3.32
Direct Kno 8.99 2.61 4.49 0.00 -0.12 0.58 7.04 3.37
Direct Goal 8.99 2.44 4.13 -0.05 -0.19 0.56 6.97 3.27
Direct Goal +Kno +Rel 9.00 2.48 3.86 -0.01 -0.13 0.68 6.85 3.25
A.4 Safety Evaluation Results
Model Benign Adversarial Benign Adversarial Benign Jailbreak
BC 0.61 0.77 0.78 0.92 0.86 0.39
Sotopia-RL 0.58 0.75 0.50 0.90 0.86 0.40
Correlation annotator1 annotator2 annotator3 annotator4 GPT-4o
annotator1 1.000 0.812 0.931 0.891 0.771
annotator2 0.812 1.000 0.737 0.717 0.636
annotator3 0.931 0.737 1.000 0.818 0.756
annotator4 0.891 0.717 0.818 1.000 0.664
GPT-4o 0.771 0.636 0.756 0.664 1.000
Table 11: Diversity evaluation results for different models. We calculate the word numbers and turn numbers for social interactions.
Metric BC Sotopia-$\pi$ GOAL-RL Sotopia-RL
Avg. Word Number 37.17 35.50 51.83 76.53
Avg. Turn Number 14.44 10.81 19.41 19.59
A.5 Diversity Evaluation Results
[…]
Appendix B Additional Analysis of Utterance-level Reward
Figure 10: Reward attribution and reward aggregation examples. On the left, it explains 4 types of attribution methods (uniform, scaled, singular, and direct). On the right, it explains 4 types of different reward aggregation methods (Rel-RL, Goal-RL, Kno-RL, Sotopia-RL) where all of them are based on direct reward attribution and Sotopia-RL is the combined one. More details are in Appendix §H and §I.
Appendix C Limitations and Future Works
Future work can extend our framework in several directions. One avenue is to explore more advanced multi-turn reinforcement learning algorithms to better capture the dynamics of long social exchanges. Another is to evaluate social agents directly in human–agent interactions, moving beyond agent–agent simulations. Finally, deploying socially intelligent agents in real-world contexts raises important challenges, including the risk of manipulative or deceptive behaviors that current testing protocols may not fully capture.
Appendix D Artifact Details
D.1 Artifact Information
@@ D.2 Artifact License
The Sotopia  [^7] https://github.com/sotopia-lab/sotopia framework used in our experiments are released under the MIT License, which permits reuse, modification, and distribution for both research and commercial purposes. The dataset we used from Sotopia-$\pi$ [^8] https://github.com/sotopia-lab/sotopia-pi are under the Apache 2.0 License. We used Sotopia and Sotopia-$\pi$ exclusively for academic research within the scope defined by this license.
D.3 Data Usage
@@ D.4 Data Statistics
LLM Annotation Data. We use GPT-4o to annotate the self-play dialogue data introduced in the “Self-Play and Behavior Cloning Data” subsection of Section D.4. The resulting annotations used for training the reward model are publicly available at https://huggingface.co/datasets/ulab-ai/sotopia-rl-reward-annotation.
Self-Play and Behavior Cloning Data. To generate GPT-4o self-play trajectories for Behavior Cloning (BC), we use a subset of Sotopia-$\pi$ consisting of 100 distinct social scenarios, each paired with two agent-specific social goals. For each scenario, GPT-4o engages in a full dialogue episode, exhibiting role-playing behaviors conditioned on these goals. We serialize these conversations and use them as training data for the BC model.
LLM Annotation Data. We use GPT-4o to annotate the self-play dialogue data introduced in the “Self-Play and Behavior Cloning Data” subsection of Section D.4.
Appendix E Experimental Details
E.1 Acronym for Experimental Settings
@@
E.2 Model Size and Budget
@@ E.5 Software Versions
Appendix F Human Evaluation Details
We provide technical details of human evaluation in this section. F.1 provides the details for human annotation system. F.2 provides the details for annotation data preparation. F.3 describes the information about human annotators. F.4 provides the details for the annotation process.
F.1 Human Annotation System
@@
Data annotation part. For the data annotation part, the annotator is guided to jump to a new page after the previously mentioned annotation instruction page. Each annotator is able to review the complete annotation example again at the data annotation page and start their official data annotation. In the data annotation part, the repeated explanation of the meaning of range for each social evaluation dimension is emphasized to make sure every annotator can understand the annotation standards correctly. Fig 13 provides an example of the instruction that annotators see for metric range explanation. Each annotator is asked to annotate the social intelligence of both agents that have a conversation. For each social intelligence dimension, annotators need to annotate the score based on the metric range and provide the reasoning for that.
@@
F.2 Annotation Data Preparation
@@
F.3 Human Annotator Information